# CoMPaSS: Enhancing Spatial Understanding in Text-to-Image Diffusion Models

### v1

@@ Abstract
Text-to-image (T2I) diffusion models excel at generating photorealistic images, but commonly struggle to render accurate spatial relationships described in text prompts. We identify two core issues underlying this common failure: 1) the ambiguous nature of spatial-related data in existing datasets, and 2) the inability of current text encoders to accurately interpret the spatial semantics of input descriptions. We address these issues with CoMPaSS, a versatile training framework that enhances spatial understanding of any T2I diffusion model. CoMPaSS solves the ambiguity of spatial-related data with the Spatial Constraints-Oriented Pairing (SCOP) data engine, which curates spatially-accurate training data through a set of principled spatial constraints. To better exploit the curated high-quality spatial priors, CoMPaSS further introduces a Token ENcoding ORdering (TENOR) module to allow better exploitation of high-quality spatial priors, effectively compensating for the shortcoming of text encoders. Extensive experiments on four popular open-weight T2I diffusion models covering both UNet- and MMDiT-based architectures demonstrate the effectiveness of CoMPaSS by setting new state-of-the-arts with substantial relative gains across well-known benchmarks on spatial relationships generation, including VISOR (+98%), T2I-CompBench Spatial (+67%), and GenEval Position (+131%). Code will be available at https://github.com/blurgyy/CoMPaSS.
@@ 1 Introduction
Recent advancements in text-to-image (T2I) diffusion models [saharia_photorealistic_2022, ramesh_hierarchical_2022, rombach_high-resolution_2022] have transformed visual content creation and significantly shaped the modern digital life [betker_improving_nodate, MidJourney, RunwayML]. These models are capable of synthesizing photorealistic images with remarkable details [dhariwal_diffusion_2021], and are now widely applied across a range of creative and practical tasks [hertz_prompt--prompt_2022, lugmayr_repaint_2022, ruiz_dreambooth_2023, poole_dreamfusion_2022]. Despite these advancements, however, a key limitation persists: today’s diffusion models frequently fail to correctly render spatial relationships described in text. For example, when given seemingly simple spatial configurations like “a motorcycle to the right of a bear”, or “a bird below a skateboard”, models that excel in realism often struggle to accurately capture these spatial relationships (Fig. 1 left).
Figure 1: CoMPaSS enables existing T2I diffusion models to generate spatial configurations that are unlikely to occur naturally and unseen during training by improving their spatial understanding.
Given that T2I diffusion models have shown significant capability in encoding abstract attributes such as aesthetic style and quality [dhariwal_diffusion_2021, rombach_high-resolution_2022, podell_sdxl_2023, esser_scaling_2024], their consistent failure in handling spatial relationships warrants closer examinations. We consider that the limitation stems from how spatial priors are organized within the training data, and how these priors are represented to the diffusion model.
@@
To better understand this limitation, we conduct a thorough investigation covering the data characteristics during training, and the language representation during generation. Our analysis identified two core issues underlying this spatial understanding deficiency. First, the organization of spatial related data in widely-used image-text datasets including LAION [schuhmann_laion-400m_2021, schuhmann_laion-5b_2022], CC-12m [changpinyo_conceptual_2021], and COCO [lin_microsoft_2014] is severely flawed. As shown in Fig. 2, spatial descriptions appear frequently ambiguous due to various reasons, such as inconsistent frame of reference, or non-spatial usage of spatial terms. This inevitably leads to the spatial terms being interpreted by the diffusion model as ambiguous. Second, commonly used text encoders including CLIP [radford_learning_2021, cherti_reproducible_2023] and T5 [raffel_exploring_2020] consistently fail to preserve spatial relationships in their embeddings. Through a proxy task where the text encoders are extensively tested to retrieve the spatial equivalent from three prompt variations (Tab. 1), we show that current text encoders struggle to interpret the accurate meanings of spatial descriptions. Furthermore, the two issues compound each other. Even if a text encoder could reliably capture spatial relationships, the lack of consistent and accurate spatial data would still limit model performance. Conversely, even high-quality spatial descriptions would be undermined by the encoder’s inability to interpret them accurately.
@@
To this end, we seek to enhance the spatial understanding of text-to-image generative diffusion models with a comprehensive solution, which we call CoMPaSS (Comprehensive Method for Positional and Spatial Synthesis). Central to our approach is the Spatial Constraints-Oriented Pairing (SCOP) data engine, which extracts object pairs with clear spatial relationships from images, along with their accurate textual descriptions. By enforcing strict criteria for visual prominence, positional clarity, and size balance, SCOP identifies and validates spatial relationships between object pairs through carefully designed spatial constraints. When applied to the COCO [lin_microsoft_2014] training split, SCOP yields a curated set of over 28,000 object pairs, each coupled with spatially-accurate text descriptions of the image content. To better exploit the accurate spatial priors, we additionally introduce Token ENcoding ORdering (TENOR), a plug-and-play module that enhances the preservation of spatial semantics from text input with explicit token ordering information, effectively compensating for the spatial understanding limitations of current text encoders. TENOR is compatible with any diffusion architecture, adds no extra parameters, and inflicts practically zero inference-time overhead.
@@
Applying CoMPaSS to popular open-weight T2I diffusion models like SD1.4, SD1.5, SD2.1 and FLUX.1 sets new state-of-the-arts with substantial relative gains on well-known benchmarks, including VISOR [gokhale_benchmarking_2023] (+98%), T2I-CompBench Spatial [huang_t2i-compbench_2023-1] (+67%), and GenEval Position [ghosh_geneval_2023] (+131%). In summary, our contributions are three-fold:
@@
  • A parameter-free module, TENOR, that further improves spatial understanding of T2I diffusion models of any architecture, in the meantime adds negligible computational overhead.
2 Related Works
Text-to-image Diffusion Models.
Diffusion models have achieved very high-quality and diverse image synthesis in recent years [ho_denoising_2020, rombach_high-resolution_2022, dhariwal_diffusion_2021, peebles_scalable_2023]. To allow natural user input through human language, text-conditional diffusion models adopt datasets with large-scale image-text pairs [changpinyo_conceptual_2021, schuhmann_laion-400m_2021, schuhmann_laion-5b_2022, liang_rich_2024], with pre-trained text encoders to bridge the gap between the two modalities. GLIDE [nichol_glide_2022], DALLE-2 [ramesh_hierarchical_2022], LDM [rombach_high-resolution_2022], and the Stable-Diffusion family of models [rombach_high-resolution_2022, podell_sdxl_2023, esser_scaling_2024] employ CLIP [radford_learning_2021, cherti_reproducible_2023] for language conditioning. PixArt-$\alpha$ [chen_pixart-alpha_2023], SD3/3.5 [esser_scaling_2024], and FLUX.1 adopt the pure-language T5 [raffel_exploring_2020] model as their text encoders. Benchmarking the alignment of image content w.r.t. the input text description has been well explored [hu_tifa_2023, ghosh_geneval_2023, gokhale_benchmarking_2023, huang_t2i-compbench_2023-1, jayasumana_rethinking_2024, hu_ella_2024]. More recently, CoMat [jiang_comat_2024] leverages image-to-text (I2T) models [li_blip_2022, wang_git_2022, liu_visual_2023] to identify the overlooked condition information in generated images during training. ParaDiffusion [wu_paragraph--image_2023] and ELLA [hu_ella_2024] explore the use of large language models (LLMs) in T2I diffusion models.
Instead of trying to enhance the language understanding ability by using more advanced text encoders, we seek an orthogonal direction of facilitating text-image alignment with a data engine that ensures accurate spatial alignment between image and text in training data.
Spatial Control in Diffusion Models.
Controllable image generation has become a popular research topic recently. Several lines of works, including both training-based and test-time optimization methods have been explored to enhance the spatial controllability of generation. Previous works add image as conditioning input [zhang_adding_2023, mou_t2i-adapter_2023, ye_ip-adapter_2023] to enable layout control, but add computational overhead to both training and inference processes. Attention-based methods build upon the relation between the text-image attention map and the layout of generated image [hertz_prompt--prompt_2022], and improve compositional generation by explicitly supervising the attention maps either during training [wang_tokencompose_2024] or at inference-time [chefer_attend-and-excite_2023, chen_training-free_2023, phung_grounded_2024]. However, supervising attention requires explicit target signal for optimization. Although works like [phung_grounded_2024, feng_layoutgpt_2024] automate the generation of target signals with LLMs, optimization on attention maps still inflicts substantial memory and speed penalty. SPRIGHT [chatterjee_getting_2024] avoids test-time optimizations by fine-tuning a diffusion model on datasets recaptioned with a vision-language model [liu_visual_2023], but requires joint training of the text encoder in addition to the diffusion network, which still adds considerable training overhead.
In contrast to existing works, we propose a network module that improves spatial understanding of T2I diffusion models of any architecture, in the meantime adds no extra parameters and inflicts practically zero computational overheads to either training or inference processes.
@@ 3 Approach
We present CoMPaSS, a comprehensive solution for improving spatial understanding in text-to-image diffusion models. In Sec. 3.1, we develop the Spatial Constraints-Oriented Pairing (SCOP) data engine, which extracts object pairs with clear spatial relationships from images, along with their accurate textual descriptions. In Sec. 3.2, we introduce the Token ENcoding ORdering (TENOR) module, a plug-and-play module that enhances the preservation of spatial semantics from text input by incorporating explicit token ordering information.
These components together form a comprehensive approach to improving spatial understanding in text-to-image diffusion models: SCOP provides high-quality spatial priors through carefully curated spatial relationships, while TENOR enables the model to effectively interpret and exploit these spatial priors during generation.
@@ 3.1 The SCOP Data Engine
High-quality spatial understanding in generative models requires training data with unambiguous spatial relationships. However, existing text-image datasets often contain problematic spatial descriptions. As shown in Fig. 2, 1) terms like “left” and “right” suffer from perspective ambiguity, referring inconsistently to viewer perspective or object-intrinsic orientation; 2) directional terms are frequently used in non-spatial contexts (e.g.the right choice”); 3) spatial relationships are often described with missing or incorrect reference objects, making the intended spatial configuration impossible to determine. These issues create significant challenges for models attempting to learn reliable spatial relationships from such data.
@@
((a)) COCO [lin_microsoft_2014]
((b)) LAION [schuhmann_laion-400m_2021, schuhmann_laion-5b_2022]
((c)) CC-12M [changpinyo_conceptual_2021]
Figure 2: Examples highlighting common ambiguities and errors in spatial language annotations from COCO, LAION, and CC-12M datasets. (a1, a2) inconsistent frame of reference; (b1) non-spatial usage of spatial terms; (b2, c1, c2) missing or incorrect reference objects.
Figure 3: Overview of the Spatial Constraints-Oriented Pairing (SCOP) data engine. SCOP first (1) reasons through all possible pairs of objects in image, then (2) identifies and validates spatial relationships between object pairs through carefully designed spatial constraints. Finally, it (3) decodes the extracted unambiguous object pair descriptors into images with accurate spatial descriptions
1. Relationship Reasoning.
The engine begins by identifying all possible pairwise relationships within each image. For every image $\mathbf{I}$, we first identify all valid object instances $\{O_{1},\cdots,O_{n}\}$, where the $i$-th object $O_{i}$ is associated with a bounding box $\mathbf{B}_{i}$ and its corresponding category label $\text{Category}(\mathbf{B}_{i})$. We then enumerate all possible pairs between the identified objects, creating a total of $(^{n}_{2})$ candidate pairs with their respective spatial relationships to evaluate.
2. Spatial Constraints Enforcement.
To ensure unambiguous spatial relationships between object pairs in images, we formalize a set of necessary constraints that collectively guarantee visual clarity and semantic validity. For any candidate object pair $(O_{i},O_{j})$, we evaluate five constraints:
@@
    A relatively small value of $\tau_{\text{v}}$ ensures the reasoned spatial relationship from last stage is a dominant feature of the image, rather than an incidental arrangement of background elements.
  • Semantic Distinction: Objects must belong to different categories, i.e.:
@@
  • Spatial Clarity: Objects must be within a reasonable proximity for clear spatial reference. Formally, let $d(i,j)=\left\lVert\text{centroid}(\mathbf{B}_{i})-\text{centroid}(\mathbf{B}_{j})\right\rVert_{2}$, and $l(i,j)=\min\left(\text{diagonal}(\mathbf{B}_{i}),\text{diagonal}(\mathbf{B}_{j})\right)$, an object pair $(O_{i},O_{j})$ is said to satisfy spatial clarity when:
@@
    With this constraint, both objects should contribute similarly to the spatial relationship, preventing cases where one object is too small to serve as a reliable spatial reference.
@@
3. Relationship Decoding.
For object pairs satisfying all constraints, we adopt a representation system that encodes the spatial relationships with structured descriptors, rather than fixed text. Such descriptors can be decoded into image-text pairs at training time. For example, a descriptor $D=$(cup, $\mathbf{B}_{\mathtt{cup}}$) <above> (couch, $\mathbf{B}_{\mathtt{couch}}$)” can be decoded into $(\mathbf{I}_{\text{roi}},\mathbf{T})$, where $\mathbf{I}_{\text{roi}}$ is a crop of the original image that contains both $\mathbf{B}_{\mathtt{cup}}$ and $\mathbf{B}_{\mathtt{couch}}$, and $\mathbf{T}$ is a text prompt formatted from a predefined pool of templates, e.g.a cup on top of a couch”, or “an image of a couch below a cup”, etc. Such a decoding process captures the exact positions of the referred objects as well as their reasoned spatial relation, allowing for both accurate relation descriptions and flexible training signals.
The SCOP Dataset.
We automate SCOP on the COCO [lin_microsoft_2014] training split to curate a dataset specifically targeted at training T2I models with improved spatial understanding capabilities, covering over 28,000 object pairs across 15,000 images, with well-defined spatial relationships and accurate spatial descriptors. This dataset significantly improves the spatial understanding of T2I diffusion models trained on it, while its size is only a fraction of web-scale datasets (0.004% of LAION-400M [schuhmann_laion-400m_2021], or 0.13% of CC-12M [changpinyo_conceptual_2021], showcasing the efficiency and efficacy of the SCOP data engine.

### v2

@@ Abstract
Text-to-image (T2I) diffusion models excel at generating photorealistic images but often fail to render accurate spatial relationships. We identify two core issues underlying this common failure: 1) the ambiguous nature of data concerning spatial relationships in existing datasets, and 2) the inability of current text encoders to accurately interpret the spatial semantics of input descriptions. We propose CoMPaSS, a versatile framework that enhances spatial understanding in T2I models. It first addresses data ambiguity with the Spatial Constraints-Oriented Pairing (SCOP) data engine, which curates spatially-accurate training data via principled constraints. To leverage these priors, CoMPaSS also introduces the Token ENcoding ORdering (TENOR) module, which preserves crucial token ordering information lost by text encoders, thereby reinforcing the prompt’s linguistic structure. Extensive experiments on four popular T2I models (UNet and MMDiT-based) show CoMPaSS sets a new state of the art on key spatial benchmarks, with substantial relative gains on VISOR (+98%), T2I-CompBench Spatial (+67%), and GenEval Position (+131%). Code is available at github.com/blurgyy/CoMPaSS.
@@ 1 Introduction
Recent advancements in text-to-image (T2I) diffusion models [62, 57, 58] have transformed visual content creation and significantly shaped the modern digital life [2, 47, 60, 4]. These models are capable of synthesizing photorealistic images with remarkable details [18], and are now widely applied across a range of creative and practical tasks [30, 45, 59, 54]. Despite these advancements, a key limitation persists: today’s diffusion models frequently fail to correctly render spatial relationships described in text. For example, when given seemingly simple spatial configurations like “a motorcycle to the right of a bear”, or “a bird below a skateboard”, models that excel in realism often struggle to accurately capture these spatial relationships. Given that T2I diffusion models have shown significant capability in encoding abstract attributes such as aesthetic style and quality [18, 58, 53, 20, 4], their consistent failure in handling spatial relationships warrants closer examination.
Figure 1: CoMPaSS enhances the spatial understanding of existing T2I diffusion models, enabling them to generate images that faithfully reflect spatial configurations specified in the text prompt.
@@
To better understand this limitation, we conduct a thorough investigation covering the essential aspects of training a diffusion model, including the characteristics of the training data and the quality of language representations used during generation. Our analysis identified two core issues underlying this spatial understanding deficiency. First, the organization of data concerning spatial relationships in widely-used image-text datasets including LAION [64, 65], CC12M [5], and COCO [43] is severely flawed. As shown in Fig. 2, spatial descriptions in these datasets are frequently ambiguous due to various reasons, such as inconsistent frames of reference or the non-spatial use of spatial terms. This inevitably leads to the spatial terms being interpreted by the diffusion model as ambiguous. Second, commonly used text encoders including CLIP [55, 14] and T5 [56] consistently fail to preserve spatial relationships in their embeddings. We demonstrate this through a proxy task (Tab. 1) where we test the ability of text encoders to identify spatially equivalent descriptions among a set of prompt variations. Our results show that these encoders struggle to interpret spatial descriptions accurately. Furthermore, the two issues compound each other. Even if a text encoder could reliably capture spatial relationships, the lack of consistent and accurate spatial data would still limit model performance. Conversely, even high-quality spatial descriptions would be undermined by the encoder’s inability to interpret them accurately.
@@
To this end, we propose CoMPaSS (Comprehensive Method for Positional and Spatial Synthesis), a comprehensive solution to enhance the spatial understanding of T2I diffusion models. Central to our approach is the Spatial Constraints-Oriented Pairing (SCOP) data engine, which extracts object pairs with clear spatial relationships from images, along with their accurate textual descriptions. By enforcing strict criteria for visual prominence, positional clarity, and size balance, SCOP identifies and validates spatial relationships between object pairs through carefully designed spatial constraints. When applied to the COCO [43] training split, SCOP yields a curated set of over 28,000 object pairs, each coupled with spatially-accurate text descriptions. To better exploit the spatial priors curated by SCOP, we introduce the Token ENcoding ORdering (TENOR) module. This plug-and-play module addresses a key limitation of text encoders by reinforcing the prompt’s linguistic structure. It does so by injecting explicit token ordering information into the model’s attention mechanism, ensuring that the sequential nature of the text is preserved. TENOR is compatible with any diffusion architecture, adds no extra parameters, and incurs negligible inference-time overhead.
@@
When applied to popular open-weight T2I diffusion models (SD1.4, SD1.5, SD2.1, and FLUX.1), CoMPaSS achieves state-of-the-art performance on well-known benchmarks, with substantial relative gains on VISOR [26] (+98%), T2I-CompBench Spatial [35] (+67%), and GenEval Position [25] (+131%). Notably, these improvements are achieved without compromising general generation capabilities, while also enhancing image fidelity. In summary, our contributions are three-fold:
@@
  • A parameter-free module, TENOR, that improves a model’s ability to interpret spatial language by preserving the crucial token order of text prompts, while adding negligible computational overhead.
2 Related Work
Text-to-image Diffusion Models. Diffusion models have achieved very high-quality and diverse image synthesis in recent years [32, 58, 18, 51, 12, 11, 10]. To enable user input via natural language, text-conditional diffusion models are trained on large-scale image-text datasets [5, 64, 65, 42], typically using pre-trained text encoders to bridge the two modalities. While early models such as GLIDE [49], DALLE-2 [57], and the Stable Diffusion family [58, 53, 20] employed CLIP [55, 14] for language conditioning, more recent work like PixArt [12, 11, 10], SD3 [20], and FLUX.1 has adopted text-only encoders like T5 [56] to further enhance faithfulness to the input prompt. To quantify this faithfulness, benchmarks with diverse evaluation methodologies have been explored [34, 25, 26, 35, 33, 39, 24, 72, 62, 83]. Although recent reinforcement learning-based methods [21, 61, 38, 69, 3, 27, 35], which often optimize for human preference using specialized datasets [42, 66, 1, 76, 75, 80, 35], have exhibited strong benchmark scores, the ambiguity underlying these datasets remains a significant challenge. Our work addresses this issue by introducing a data engine designed to ensure accurate spatial alignment between images and text in the training data.
Spatial Control in Diffusion Models. Spatial control is one of the most popular demands in image generation [28, 63, 85, 29, 77, 50]. Approaches for controllable text-to-image generation can be broadly categorized by whether they manipulate the training or the inference stage of a diffusion model. Training-based methods [59, 84, 48, 81, 6, 71] adapt the pretrained weights of diffusion models for specific downstream tasks [78, 40, 70, 82, 73, 71, 6, 33]. However, they generally incur significant training overhead due to architectural augmentations [84, 33, 74], extra training modules [84, 6], or the need for additional supervision [71, 37]. Inference-only methods [13, 79, 52, 44, 22, 7, 30, 67, 46] achieve robust targeted generation and editing, but often rely on explicit inputs like bounding boxes. These inputs may require manual annotation or specialist models that have separate training costs [23, 15] or high inference costs [41, 79]. Furthermore, many of these approaches introduce significant inference overhead due to costly test-time optimization. In contrast, our work introduces a two-part solution. Our SCOP data engine provides high-quality training data, while our TENOR module ensures the model can leverage this data by preserving the crucial token order of text prompts. This combined approach efficiently improves spatial understanding without extra trainable parameters or significant computational overhead.
@@ 3 Approach
We present CoMPaSS, a comprehensive solution for improving spatial understanding in text-to-image diffusion models. In Sec. 3.1, we introduce the Spatial Constraints-Oriented Pairing (SCOP) data engine, which extracts object pairs with clear spatial relationships from images, along with their accurate textual descriptions. In Sec. 3.2, we introduce the Token ENcoding ORdering (TENOR) module, a plug-and-play module that improves the model’s ability to process spatial language by preserving the crucial token order of text prompts.
These two components work in synergy: SCOP provides high-quality spatial priors through carefully curated spatial relationships, while TENOR ensures the model can distinguish between structurally different prompts, allowing it to better leverage these spatial priors during generation.
@@ 3.1 The SCOP Data Engine
High-quality spatial understanding in generative models requires training data with unambiguous spatial relationships. However, existing text-image datasets often contain problematic spatial descriptions. As shown in Fig. 2, 1) terms like “left” and “right” suffer from perspective ambiguity, referring inconsistently to viewer perspective or object-intrinsic orientation; 2) directional terms are frequently used in non-spatial contexts (e.g. “the right choice”); 3) spatial relationships are often described with missing or incorrect reference objects, making the intended spatial configuration impossible to determine. These issues create significant challenges for models attempting to learn reliable spatial relationships from such data.
@@
((a)) COCO [43]
((b)) LAION [64, 65]
((c)) CC12M [5]
Figure 2: Examples highlighting common ambiguities and errors in spatial language annotations from COCO, LAION, and CC12M datasets. (a1, a2) inconsistent frame of reference; (b1) non-spatial usage of spatial terms; (b2, c1, c2) missing or incorrect reference objects.
Figure 3: Overview of the Spatial Constraints-Oriented Pairing (SCOP) data engine. SCOP first (1) reasons about all possible object pairs in an image, then (2) validates their spatial relationships using a set of carefully designed constraints. Finally, it (3) decodes the resulting unambiguous descriptors into image crops paired with accurate textual descriptions.
1. Relationship Reasoning. The engine begins by identifying all possible pairwise relationships within each image. For every image $\mathbf{I}$, we first identify all valid object instances $\{O_{1},\cdots,O_{n}\}$, where the $i$-th object $O_{i}$ is associated with a bounding box $\mathbf{B}_{i}$ and its corresponding category label $\text{Category}(\mathbf{B}_{i})$. We then enumerate all possible pairs of identified objects, creating a total of $\binom{n}{2}$ candidate pairs to be evaluated for their spatial relationships.
2. Spatial Constraints Enforcement. To ensure unambiguous spatial relationships between object pairs in images, we formalize a set of necessary constraints that collectively guarantee visual clarity and semantic validity. For any candidate object pair $(O_{i},O_{j})$, we evaluate five constraints:
@@
    A relatively small value of $\tau_{\text{v}}$ ensures that the identified spatial relationship is a dominant feature of the image, rather than an incidental arrangement of background elements.
  • Semantic Distinction: Objects must belong to different categories, satisfying the following condition:
@@
  • Spatial Clarity: Objects must be within a reasonable proximity for clear spatial reference. Formally, we define the distance between centroids as $d(i,j)=\left\lVert\text{centroid}(\mathbf{B}_{i})-\text{centroid}(\mathbf{B}_{j})\right\rVert_{2}$ and a characteristic length as $l(i,j)=\min\left(\text{diagonal}(\mathbf{B}_{i}),\text{diagonal}(\mathbf{B}_{j})\right)$. An object pair $(O_{i},O_{j})$ satisfies this constraint when:
@@
    This constraint ensures that both objects contribute similarly to the spatial relationship, preventing cases where one object is too small to serve as a reliable spatial reference.
@@
3. Relationship Decoding. For object pairs satisfying all constraints, we adopt a representation system that encodes the spatial relationships with structured descriptors, rather than fixed text. These descriptors are then decoded into image-text pairs at training time. For example, a descriptor such as D=D=(cup, $\mathbf{B}_{\mathtt{cup}}$) <above> (couch, $\mathbf{B}_{\mathtt{couch}}$)” is decoded into a pair $(\mathbf{I}_{\text{roi}},\mathbf{T})$. Here, $\mathbf{I}_{\text{roi}}$ is a crop of the original image containing both $\mathbf{B}_{\mathtt{cup}}$ and $\mathbf{B}_{\mathtt{couch}}$, and $\mathbf{T}$ is a text prompt generated from a predefined pool of templates (e.g., “a cup on top of a couch” or “an image of a couch below a cup”). This decoding process captures the exact positions of the referenced objects and their identified spatial relationship, providing both accurate textual descriptions and flexible training signals.
The SCOP Dataset. We apply the SCOP engine to the COCO [43] training split to curate a dataset for improving the spatial understanding of T2I models. The resulting dataset contains over 28,000 object pairs from 15,000 images, each characterized by a well-defined spatial relationship and an accurate textual descriptor. The symbolic descriptors used during SCOP’s curation process are internal to the data engine and are not used at inference time. The final trained model takes only standard natural language prompts as input. The small size of this dataset, which is only a fraction of web-scale datasets like LAION-400M (0.004%) [64] and CC12M (0.13%) [5], highlights the data efficiency of our targeted curation approach. To validate the quality of our curation process, we conduct a human study following VISOR’s evaluation protocol, finding an 85.2% agreement rate between human annotations and our generated image-text pairs, indicating strong alignment.